OpenStudy (anonymous):

Solve the equation on the interval [0,2pi): 12cos^2x-9=0

OpenStudy (campbell_st):

well start by rewriting the equation \[12\cos^2(x) = 9\] divide both sides by 12 \[\cos^2(x) = \frac{9}{12}\] simplify the fraction and then take the square root of both sides... this will result in 2 answers for cos(x) positive and negative. Which means you will have 4 answers as answers on your given domain.

OpenStudy (anonymous):

Thanks for the help I got my 4 answers.